OpenStudy (anonymous):

guys....expand dis: (a+b)^3.....n a^3 + b^3

OpenStudy (anonymous):

(a+b)^3 = (a+b)(a+b)(a+b) = (a^2+2ab+b^)(a+b) = a^3+3(a^2)b+3a(b^2)+b^3

OpenStudy (anonymous):

sorry the third line is supposed to say (a^2+2ab+b^2)(a+b)

OpenStudy (anonymous):

thnx:)